# Guiding Vision-Language Model Selection for Visual Question-Answering Across Tasks, Domains, and Knowledge Types

### v1

@@ 4 Comparative Evaluation of VLMs
We evaluate state-of-the-art (SOTA) VLMs in practical settings using our framework. We use GoEval-mini for all the evaluations, since it demonstrated maximum correlation with human judgements, and is more cost-efficient.
For the VLMs, we use InternVL-2 1B and 8B [10, 9], PaliGemma-3B [4], Qwen-2-VL 2B and 7B [2, 34], LlaVa-1.6-Mistral-7B [22], CogVLM2-Llama-3-19B [36, 15], Gemini-1.5 Flash and Pro [33], and GPT-4o-Mini [30]. The rationale behind choosing these models is to have sufficient diversity to allow users to choose appropriate VLM based on other constraints.
All models except Gemini-1.5 and GPT-4o are open sourced, which gives freedom to customize the models as desired, and host it in-house. It takes away the privacy and regulatory risk of sending data to a third-party. It also reduces operational and opportunity cost factor, as these APIs have high associated cost and rate limits. In-house hosting allows a relatively fixed cost rather than a variable cost with usage. Out of these open-sourced VLMs, we have taken smaller models in 1B-3B range too, which can be used in resource-constrained environments and on-device AI, if desired. Larger models may give better performance, but for more resource. If an application expects SOTA performance, doesn’t have constraints of API cost or rate limits, and doesn’t want any constraints of hosting the models, Gemini-1.5 and GPT-4o models can be utilized directly.
For all the open-sourced models, we use HuggingFace implementation with the image and question in a prompt recommended by the model card, since we want to evaluate all scenarios unformly. For Gemini models, we use the Gemini APIs444https://ai.google.dev/gemini-api/docs, and for GPT-4o-mini, we use the OpenAI API555https://platform.openai.com/docs/overview. The results are discussed in the following subsections. More details about the artifacts used are given in Table 5 of the Appendix A.1 (in supplementary material).
@@ 4.1 Correlation Between VLM Outputs
Our first hypothesis was that all VLMs don’t perform similar with all task instances. They have their own strengths and weaknesses. To establish that, we evaluate the correlation between outputs of different models for all task instances, and document the results in Figure 4.
From the figure, most of the correlations are blue, resembling low correlation between outputs. This shows that the performance difference between VLMs is not just in terms of a global statistic, but also differs for individual task instances. Thus, some VLMs that might be great on a subset of tasks, may be poor on another subset.
The highest correlation is observed between Gemini-1.5 Flash and Pro, followed by Qwen-2-VL 2B and 7B, and InternVL-2-1B and 8B. This behavior can be explained by the fact that these VLMs are part of the same family. They might have similar training data, training strategy, and architecture. Some open-source VLMs also exhibit higher correlations (light blue) with the three proprietary models—Gemini-1.5-Pro, Gemini-1.5-Flash, and GPT-4o-mini—though these correlations remain relatively low.
Since task types, application domains and knowledge types are key factors on which tasks can be differentiated in practical settings, we move to analyzing the performance on those factors in the following subsections.
@@ 4.2 Evaluation on Task Types
We evaluated VLMs on four different task types - Chart Understanding using ChartQA dataset [26], Document Understanding using DocVQA dataset [27], Knowledge-based VQA using A-OKVQA [31] and OK-VQA [25], and general VQA using VQAv2 [13]. The performance of different VLMs across these task types is summarized in Table 4.
@@
Chart Understanding | 46.81 | 12.31 | 12.84 | 22.88 | 14.06 | 65.68 | 67.34 | 62.17 | 67.66 | 60.36 |
Doc. Understanding | 45.68 | 15.81 | 12.49 | 30.07 | 20.00 | 65.76 | 68.17 | 64.36 | 70.53 | 54.68 |
Knowledge-based VQA | 45.41 | 19.83 | 19.22 | 58.65 | 26.94 | 71.43 | 62.77 | 70.66 | 77.31 | 84.04 |
@@
From the table, the closed models, which are believed to be SOTA, outperform the smaller, open-sourced models. This is expected considering those models are larger. Within them, we identify a pattern that Gemini-1.5-Pro performs significantly better than GPT-4o-mini when visual inference and extracting information from image is more important, like interpreting documents and charts. On the other hand, where knowledge and comprehension is critical with tricky questions using standard images, GPT-4o-mini outperforms Gemini-1.5-Pro. Therefore, based on the nature of the problem, one of these models can be chosen. If cost of processing is a factor, Gemini-1.5-Flash can be chosen with approximately 7-10% performance loss.
Among the open-sourced models, InternVL-2 shows promising results with both 1B and 8B models given their size, and can be chosen if open-sourced models are needed for their advantages, based on resource availability. CogVLM-2-Llama-3-19B also competes closely with Gemini-1.5-Flash in Chart and Document understanding tasks. Llava-1.6-Mistral-7B performs acceptable in knowledge-based VQA and VQA, but the performance degrades drastically in the other two categories, where visual abilities are critical, exposing its limitations in that area. Qwen-2-VL variants and PaliGemma-3B surprisingly prove to be weak in all tasks.
@@ 4.3 Evaluation on Application Domains
We evaluate the VLMs on all application domains which had more than 300 task instances, and the results are shown in Figure 5.
If the VLMs would be similar in all application domains, their result will be perfectly circular. But, we see that most of the VLM performance graphs are not circular and have aberrations in multiple categories, which establish that VLMs have weaknesses and strengths. In most cases, the weakness of one of few VLMs is compensated in the strength of others. So, choosing a VLM wisely as per the application need can help mitigating the weaknesses.
Among the closed models, GPT-4o-mini proves to be the best in four categories - Nature, Nutrition and Food, Social Media and Sports, and Gemini-1.5-Pro proves to be the best in all other categories. GPT-4o-mini doesn’t even remain second best in some categories like Mathematics, Economics, Law, and is outperformed by many open-sourced models as well in these categories. Thus, if one uses this model in these, they may not achieve results even close to best possible. Gemini-1.5-Flash remains strong with a performance deficit compared to Pro, but in Social Media tasks, it almost matches Pro. Therefore, while appropriate model should be selected based on requirement, and Gemini-1.5-Pro generally looks to be the best overall choice.
In the open-sourced model category, InternVL-2-8B and CogVLM-2-Llama-3-19B are the best possible choices. CogVLM-2-Llama-3-19B generally performs well in more academic topics like Mathematics, Computer Science, law, Government and Politics, but suffers a lot of performance degradation in more social topics like Nature, Nutrition and Food, Social Media, Sports. So, it may be a decent choice for the earlier, but not for the latter categories. InternVL-2-8B also shows similar traits, but the difference is relatively less. For academic topics, these models prove particularly strong, sometimes even outperforming some of the the closed models. Llava-1.6-Mistral-7B is one model that shows exactly opposite trait than this, being limited in academic topics as compared to social topics, but its performance remains at a gap compared to the two previously mentioned models. Qwen-2 variants and PaliGemma show weak results in all domains, like in task types. InternVL-2-1B remains the best choice if small-sized model is required, with decent results using 1B parameters.
@@ 4.4 Evaluation on Knowledge Types
Similar to application domains, we evaluate all VLMs on all knowledge types where number of task instances are greater than 300. We demonstrate the results using a similar radar chart in Figure 6.
In the closed models category, the SOTA performance is again shared by Gemini-1.5-Pro and GPT-4o-mini. But unlike in application domains, GPT-4o-mini competes more strongly with Gemini-1.5-Pro on several knowledge types. Visual, Social and Commonsense knowledge is where the advantage of GPT-4o-mini over Gemini-1.5-Pro is maximum, again depicting its strength in social knowledge types. It’s still considerably weak in temporal, scientific, and mathematical knowledge, falling behind even some of the open-sourced small models. Gemini-1.5-Pro will still be the better choice overall, but different models can be selected based on the results obtained for specific knowledge types. Gemini-1.5-Flash again follows similar trend, with slight performance difference from Pro.
Comparing the open-sourced models, InternVL-2-8B and CogVLM-2-Llama-3-8B are the best choices here as well, but here we can see that InternVL-2-8B outperforms the latter in all knowledge types. InternVL-2-1B continues to be the best overall choice if small models are required, and Qwen-2 variants and PaliGemma continue to prove to be least effective. Therefore, InternVL-2-8B is the best overall choice for open-sourced models.
@@ 4.5 Overall Analysis and Recommendations
In subsections 4.2, 4.3 and 4.4, we evaluated all vision-language models under three different aspects of task types, application domains and knowledge types. We also discussed their strengths and weaknesses in those categories and recommended different VLMs to use under different requirements. In this section, we will take a higher-level look considering everything together.
We identified that Gemini-1.5-Pro and GPT-4o-mini, both are strong models. In general, GPT-4o-mini is found to be weak in analytical tasks, which belonged to Mathematics or Economics domain, or scientific or mathematical knowledge types. It not only falls behind Gemini-1.5-Pro in these tasks, but also behind open models like InternVL-2-8B or CogVLM-2-Llama-3-19B. However, it is strong is social and topical tasks like Nature, Sports, Nutrition and Food domains. Gemini-1.5-Pro is a strong model across multiple usages. It is either the best, or comes close second or third best. However, this model is expensive to use. Therefore, if cost is a significant factor, Gemini-1.5-Flash can be considered as a decent alternative, which performs in similar trends as the Pro, but at a performance deficit of around 7-10%. Despite this difference, its overall performance is very strong, considering that it’s significantly cost efficient in comparison to Pro.
InternVL-2-8B and CogVLM-2-Llama-3-19B are strong open-sourced models. Due to size differences, resource availability also contributes in deciding which model to use. CogVLM-2-Llama-3-19B is better at more academic tasks, that belong to domains like History, Law, Computer Science, etc., or knowledge types like Temporal, Scientific or Mathematical Knowledge. InternVL-2-8B is a more overall capable model, proving to be more suited in broader application requirements. It is generally the best or close second-best in in open-sourced models, and also outperforms Gemini-1.5-Flash in many categories. In some cases, it outperforms GPT-4o-mini as well. Possessing superior performance in addition to other advantages of open-sourced model makes it a strong choice. One should consider it if open-souced models are required and sufficient resources are available to host the model. It can also be aligned for downstream tasks to improve performance.
@@
The Qwen-2-VL variants and PaliGemma-3B didn’t prove fit for use in our experimental settings, being very weak on all categories. LlaVa-1.6-Mistral-7B also performs average, similar to InternVL-2-1B, but is weak in all aspects compared to InternVL-2-8B, a similar sized model.
Since Gemini-1.5-Pro was the most successful model, we demonstrate some of the qualitative examples using that model in Table 9 in the Appendix A (in supplementary material). At last, Table 10 and 11 of the Appendix A (in the supplementary material) contain quantitative results on all application domains and knowledge types respectively, including the categories that were excluded from the study of the main paper.
@@ 5 Conclusion
In this paper, we propose a comprehensive framework for evaluating Vision-Language Models (VLMs) across a wide spectrum of visual question-answering (VQA) tasks. The framework enables a rigorous comparison of VLMs based on specific application requirements. We introduce a novel evaluation paradigm that classifies VQA tasks according to three key dimensions: task types, application domains, and knowledge types. To facilitate this, we release a dataset derived from established VQA benchmarks, and annotated 4 four task types, 22 application domains, and 15 knowledge types for classification.
Additionally, we present a new evaluation metric, GoEval, which demonstrates a correlation factor of 56.71% with human judgments, making it a superior choice compared to traditional metrics. GoEval leverages the capabilities of GPT-4o for evaluation, incorporating both visual and textual information to deliver more nuanced performance assessments.
Through extensive experimentation with 10 state-of-the-art VLMs, we find that model performance varies significantly across different categories, with no single VLM emerging as universally optimal. Notably, proprietary models such as Gemini-1.5-Pro and GPT-4o-mini achieve the highest performance overall, while open-source models like InternVL-2-8B and CogVLM-2-Llama-3-19B, exhibit competitive results in specific scenarios. Our findings offer practical insights for selecting VLMs based on task-specific requirements and resource constraints, establishing a standardized framework for VLM evaluation and selection in downstream tasks. Furthermore, this evaluation framework can be extended to other vision-language tasks beyond VQA, fostering advancements in multimodal research.
@@ Appendix A Appendix
In this appendix, we provide additional details and results related to our work. Implementation details can be found in Section A.1. Table 7 provides prompts designed to classify tasks by domain (e.g., "Anthropology," "Computer Science") and by the type of knowledge required (e.g., "Commonsense Knowledge," "Visual Knowledge"). Table 8 presents the prompts used in the GoEval to verify whether a candidate answer is correct, with or without reference answers, both for visual and text-only evaluation. Additionally, Table 9 details some qualitative examples using the best overall model that we found - Gemini-1.5-Pro. At last, Table 10 and 11 contain quantitative results on all application domains and knowledge types respectively, including the categories that were excluded from the study of the main paper.
@@ A.1 Implementation Details
In this subsection, we will discuss more around implementation details. Table 5 contains all the model cards, which contain exact details about how we implemented all the VLMs, and the recommended prompt templates that we used in our evaluation.
Artifact | Implementation Link |
@@
For all the models, we prepend the question with a static text saying ‘Only answer the below question. Do not provide any additional information’. Further, we resize all the images to before sending through the model. This is the input to every model. For decoding the output, we use Greedy sampling, since fluency is not a key factor in VQA as long as the answers are correct. We use for all models.
The implementation of executing all models and our evaluation metric can be found in the Code provided. They will be shared separately as a repository with the camera ready version, along with the dataset. We also provide the implementation of generating captions and object tags, which can be used if this framework is being adapted to other tasks. All the configuration parameters and hardware used are detailed in Table 6.
@@
We also use 4-bit quantization and Flash Attention 2 [11] wherever supported for memory and execution efficiency.

### v3

@@ 4 Comparative Evaluation of VLMs
We evaluate state-of-the-art (SOTA) VLMs in practical settings using our framework using GoEval-mini, since it demonstrated maximum correlation with human judgments, and is more cost-efficient.
For VLMs, we use InternVL-2 1B and 8B Chen et al. (2023, 2024), PaliGemma-3B Beyer* et al. (2024), Qwen-2-VL 2B and 7B Bai et al. (2023); team (2024), LlaVa-1.6-Mistral-7B Liu et al. (2024), CogVLM2-Llama-3-19B Wang et al. (2023); Hong et al. (2024), Gemini-1.5 Flash and Pro Team et al. (2024), and GPT-4o-Mini OpenAI et al. (2024). The rationale behind choosing these models is to have sufficient diversity to allow users to choose the appropriate VLM based on other constraints.
All models except Gemini-1.5 and GPT-4o are open sourced, which gives freedom to customize the models as desired, and host it in-house. It takes away the privacy and regulatory risk of sending data to a third-party, and reduces operational and opportunity cost factor, as these APIs are costly with rate limits. In-house hosting allows a relatively fixed cost. We have taken smaller models in the 1B-3B range which can be used in resource-constrained environments and on-device AI.
For all open-source models, we use the HuggingFace implementation with the image and question in a prompt recommended by the model card, since we want to evaluate all scenarios uniformly. We use Gemini APIs333https://ai.google.dev/gemini-api/docs OpenAI API444https://platform.openai.com/docs/overview for Gemini and GPT-4o-mini. The results are discussed in the following subsections. More details of the artifacts used are given in Table 5 of the Appendix A.1.
@@ 4.1 Correlation Between VLM Outputs
Our first hypothesis was that all VLMs do not perform similarly with all task instances. They have their own strengths and weaknesses. To establish that, we evaluated the correlation between the outputs of different models for all tasks and document the results in Figure 4.
From the figure, most of the correlations are low. This shows that the performance difference between VLMs is not just in terms of a global statistic, but also differs for individual task instances. Thus, some VLMs that might be great in a subset of tasks may be poor in another subset.
The highest correlation is observed between Gemini-1.5 Flash and Pro, followed by Qwen-2-VL 2B and 7B, and InternVL-2-1B and 8B. Since they are from the same family, they might have similar training data, training strategy, and architecture. Some open-source VLMs also exhibit higher correlations (light blue) with Gemini and GPT, though these correlations remain relatively low.
Since task types, application domains, and knowledge types are key factors in which tasks can be differentiated in practical settings, we move to analyzing the performance on those factors.
@@ 4.2 Evaluation on Task Types
We evaluated VLMs on four different task types: chart understanding using ChartQA dataset Masry et al. (2022), Document Understanding using DocVQA dataset Mathew et al. (2021), knowledge-based VQA using A-OKVQA Schwenk et al. (2022) and OK-VQA Marino et al. (2019), and general VQA using VQAv2 Goyal et al. (2017a). The performance of VLMs across these is summarized in Table 4.
@@
Chart U. | 46.81 | 12.31 | 12.84 | 22.88 | 14.06 | 65.68 | 67.34 | 62.17 | 67.66 | 60.36 |
Document U. | 45.68 | 15.81 | 12.49 | 30.07 | 20.00 | 65.76 | 68.17 | 64.36 | 70.53 | 54.68 |
KBQA | 45.41 | 19.83 | 19.22 | 58.65 | 26.94 | 71.43 | 62.77 | 70.66 | 77.31 | 84.04 |
@@
From the table, the closed models, which are believed to be SOTA, outperform the smaller, open-sourced models. This is expected considering that those models are larger. Within them, we identify that Gemini-1.5-Pro performs significantly better than GPT-4o-mini when extracting information from an image is key, like interpreting documents and charts. On the other hand, where knowledge and comprehension are critical, GPT-4o-mini outperforms Gemini-1.5-Pro. If processing cost is a factor, Gemini-1.5-Flash can be chosen with approximately 7-10% performance loss.
Among open-source models, InternVL-2 shows promising results with both 1B and 8B models given their size, and can be chosen if open-source models are needed. CogVLM-2-Llama-3-19B also competes closely with Gemini-1.5-Flash in Chart and Document understanding tasks. Llava-1.6-Mistral-7B performs acceptable in knowledge-based VQA and VQA, but the performance degrades drastically in the other two categories, where visual comprehension is critical, exposing its limitations in that area. Qwen-2-VL variants and PaliGemma-3B surprisingly prove to be weak in all tasks.
@@ 4.3 Evaluation on Application Domains
We evaluated the VLMs in all application domains that had more than 300 task instances, and the results are shown in Figure 5.
If the VLMs were similar in all application domains, their result would be perfectly circular. However, we see that most of the VLM performance graphs have aberrations in multiple categories, highlighting variance in strengths. In most cases, the weakness of one of the few VLMs is compensated for by the strength of others. So, choosing a VLM wisely according to the application need can help mitigate weaknesses.
Among the closed models, GPT-4o-mini proves to be the best in four categories - Nature, Nutrition and Food, Social Media and Sports, and Gemini-1.5-Pro proves to be the best in all other categories. GPT-4o-mini doesn’t even remain second best in some categories like Mathematics, Economics, Law, and is outperformed by many open-sourced models here. Gemini-1.5-Flash remains strong with a performance deficit compared to Pro, but in Social Media tasks, it almost matches Pro. Therefore, while the appropriate model should be selected based on requirements, the Gemini-1.5-Pro generally looks to be the best overall choice.
In the open-source model category, InternVL-2-8B and CogVLM-2-Llama-3-19B are the best possible choices. CogVLM-2-Llama-3-19B generally performs well in more academic topics like Mathematics, Computer Science, law, Government and Politics, but suffers a lot of performance degradation in more social topics like Nature, Nutrition and Food, Social Media, Sports. InternVL-2-8B also shows similar traits, but the difference is relatively less. For academic topics, these models even outperform some of the closed models. Llava-1.6-Mistral-7B is one model that shows exactly opposite trait than this, being limited in academic topics as compared to social topics. Qwen-2 variants and PaliGemma show weak results in all domains, like in task types. InternVL-2-1B remains the best choice if a small model is required, with decent results using 1B parameters.
@@ 4.4 Evaluation on Knowledge Types
Similar to application domains, we evaluate all VLMs on all knowledge types where number of task instances is greater than 300. We demonstrate the results using a similar radar chart in Figure 6.
In the closed models category, the SOTA performance is again shared by Gemini-1.5-Pro and GPT-4o-mini. But unlike in application domains, GPT-4o-mini competes more strongly with Gemini-1.5-Pro on several knowledge types. Visual, Social and Commonsense knowledge is where the advantage of GPT-4o-mini over Gemini-1.5-Pro is maximum, again depicting its strength in social knowledge types. It is still considerably weak in temporal, scientific, and mathematical knowledge, falling behind even some of the small open-source models. Gemini-1.5-Pro will still be the best overall choice, but different models can be selected based on the results obtained for specific knowledge types. Gemini-1.5-Flash again follows similar trend, with slight performance difference from Pro.
Comparing open-source models, InternVL-2-8B and CogVLM-2-Llama-3-8B are the best choices here as well, but here we can see that InternVL-2-8B outperforms the latter in all knowledge types. InternVL-2-1B continues to be the best overall choice if small models are required, and the Qwen-2 variants and PaliGemma continue to be the least effective. Therefore, InternVL-2-8B is the best overall choice for open source models.
@@ 4.5 Overall Analysis and Recommendations
In subsections 4.2, 4.3 and 4.4, we evaluated all vision-language models under three different aspects of task types, application domains and knowledge types. We also discussed their strengths and weaknesses in those categories and recommended different VLMs to use under different requirements. In this section, we will take a higher level look at everything together.
We identified that Gemini-1.5-Pro and GPT-4o-mini are different. In general, GPT-4o-mini is weaker in analytical tasks, like tasks of Mathematics or Economics domain, or scientific or mathematical knowledge types. It not only falls behind Gemini-1.5-Pro in these tasks, but also behind open models like InternVL-2-8B or CogVLM-2-Llama-3-19B. However, it is strong is social and topical tasks like Nature, Sports, Nutrition and Food domains. Gemini-1.5-Pro proves strong more generally, but is expensive. It is either the best, or comes close second or third best. Therefore, if cost is a significant factor, Gemini-1.5-Flash can be considered as a decent alternative at a performance deficit of around 7-10%.
InternVL-2-8B and CogVLM-2-Llama-3-19B are strong open-source models. Due to size differences, resource availability also contributes to deciding which model to use. CogVLM-2-Llama-3-19B is better at more academic tasks, that belong to domains like History, Law, Computer Science, etc., or knowledge types like Temporal, Scientific or Mathematical Knowledge. InternVL-2-8B is a more general capable model that demonstrates more suitability in broader application requirements. In some cases, it outperforms GPT-4o-mini as well. Possessing superior performance in addition to other advantages of open-sourced model makes it a strong choice. It can also be aligned for downstream tasks to improve performance.
@@
The Qwen-2-VL variants and PaliGemma-3B did not prove fit for use in our experimental settings, being very weak on all categories. LlaVa-1.6-Mistral-7B also performs average, similar to InternVL-2-1B, but is weak in all aspects compared to InternVL-2-8B, a similar-sized model.
Since Gemini-1.5-Pro was the most successful model, we demonstrate some of the qualitative examples using that model in Table 9 in the Appendix A. Finally, Tables 10 and 11 of the Appendix A contain quantitative results on all application domains and knowledge types, respectively, including the categories that were excluded from the study of the main paper.
@@ 5 Conclusion
In this paper, we propose a comprehensive framework for evaluating Vision-Language Models (VLMs) across diverse visual question-answering (VQA) tasks, addressing specific application requirements. Our framework introduces a novel evaluation paradigm that classifies VQA tasks along three dimensions: task types, application domains, and knowledge types. To support this, we release VQA360, a dataset annotated across 4 task types, 22 application domains, and 15 knowledge types, derived from established VQA benchmarks. We also present GoEval, a new evaluation metric to complement it, leveraging GPT-4o to integrate visual and textual information, achieving a 56.71% correlation with human judgments and outperforming traditional metrics.
Through experiments with 10 state-of-the-art VLMs, we observe significant performance variation across categories, with no single model proving universally optimal. Proprietary models like Gemini-1.5-Pro and GPT-4o-mini achieve the highest overall performance, while open-source models such as InternVL-2-8B and CogVLM-2-Llama-3-19B excel in specific scenarios. Our findings provide actionable insights for task-specific VLM selection, and establishes a evaluation framework that can be extended to other vision-language tasks, fostering progress in multimodal research.
@@ Appendix A Appendix
In this appendix, we provide additional details and results related to our work. The implementation details can be found in Section A.1. Table 7 provides prompts designed to classify tasks by domain (e.g., "Anthropology," "Computer Science") and by the type of knowledge required (e.g., "Commonsense Knowledge," "Visual Knowledge"). Table 8 presents the prompts used in GoEval to verify whether a candidate answer is correct, with or without reference answers, both for visual and text-only evaluation. Additionally, Table 9 details some qualitative examples using the best overall model that we found - Gemini-1.5-Pro. Finally, Tables 10 and 11 contain quantitative results on all application domains and knowledge types, respectively, including the categories that were excluded from the study of the main paper.
@@ A.1 Implementation Details
In this subsection, we will discuss more details around implementation. Table 5 contains all the model cards, which contain exact details of how we implemented all the VLMs, and the recommended prompt templates that we used in our evaluation.
Artifact | Link |
@@
For all the models, we prepend the question with a static text saying ‘Only answer the below question. Do not provide any additional information. In addition, we resize all images to before sending them through the model. This is the input to every model. For decoding the output, we use Greedy sampling, since fluency is not a key factor in VQA as long as the answers are correct. We use for all models.
The implementation of executing all models and our evaluation metric can be found in the Code provided. We also provide the implementation of captions and object tags that can be used if this framework is being adapted to other tasks. All the configuration parameters and hardware used are detailed in Table 6.
@@
We also use 4-bit quantization and Flash Attention 2 Dao (2023) wherever supported for memory and execution efficiency.